# DBLP: Noise Bridge Consistency Distillation For Efficient And Reliable Adversarial Purification

### v1

@@ Introduction
0: Dataset $\mathcal{D}$, LCM $\boldsymbol{f_{\theta}}$ and its initial model parameter $\boldsymbol{\theta}$, classifier $\mathrm{C}$, ground truth label $y_{\mathrm{true}}$, Leapfrog ODE solver $\Psi_{\mathrm{leap}}$, distance metric $d(\cdot,\cdot)$, EMA rate $\mu$, noise schedule $\alpha_{t}$, skip interval $k$, encoder $\mathcal{E}$; $\boldsymbol{\theta}^{-}\leftarrow\boldsymbol{\theta}$;
@@
4:  $\boldsymbol{\epsilon}_{a}=\arg\max_{\epsilon}\mathcal{L}(\mathrm{C}(\mathcal{D}(\mathbf{z}+\boldsymbol{\epsilon})),y_{\mathrm{true}})$;
5:  $\tilde{\mathbf{z}}_{t_{n+k}}$
6:  $=\sqrt{\bar{\alpha}_{t_{n+k}}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t_{n+k}}}\boldsymbol{\epsilon}+\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t_{n+k}})}{\sqrt{\bar{\alpha}_{t_{n+k}}}(1-\bar{\alpha}_{T})}\boldsymbol{\epsilon}_{a}$;
7:  ${\hat{\mathbf{z}}_{t_{n}}}^{\Psi_{\mathrm{leap}}}\leftarrow\mathbf{z}_{t_{n+k}}+\Psi(\mathbf{z}_{t_{n+k}},t_{n+k},t_{n},\varnothing)$;
8:  $\mathcal{L}_{\mathrm{CD}}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{-}\right)$
@@
9:  $=d\left(\boldsymbol{f_{\theta}}\left(\tilde{\mathbf{z}}_{t_{n+k}},\varnothing,t_{n+k}\right),\boldsymbol{f_{\theta^{-}}}\left({\hat{\mathbf{z}}_{t_{n}}}^{\Psi_{\mathrm{leap}}},\varnothing,t_{n}\right)\right)$;
@@
10:  $\mathcal{L}_{\mathrm{rec}}\left(\boldsymbol{\theta}\right)=d\left(\boldsymbol{f_{\theta}}\left(\tilde{\mathbf{z}}_{t},\varnothing,t\right),\mathbf{z}\right)$;
11:  $\mathcal{L}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{-}\right)=\mathcal{L}_{\mathrm{CD}}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{-}\right)+\lambda_{\mathrm{rec}}\mathcal{L}_{\mathrm{rec}}\left(\boldsymbol{\theta}\right)$;
12:  $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\theta}\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\theta}^{-})$;
13:  $\boldsymbol{\theta}^{-}\leftarrow\texttt{sg}(\mu\boldsymbol{\theta}^{-}+(1-\mu)\boldsymbol{\theta})$
14:end while
@@
Following LCM (Luo et al. 2023a), let $\mathcal{E}$ and $\mathcal{D}$ denote the encoder and decoder that map images to and from the latent space, respectively. Given an image $\mathbf{x}$, its latent representation is $\mathbf{z}=\mathcal{E}(\mathbf{x})$. Unlike DDPM, DBLP includes adversarial perturbations $\boldsymbol{\epsilon}_{a}$ at the start of the noising process, such that $\mathbf{z}_{0}^{a}=\mathbf{z}_{0}+\boldsymbol{\epsilon}_{a}$. The perturbed forward process then becomes $\mathbf{x}_{t}^{a}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@
Our objective is to learn a trajectory that maps the diffused adversarial distribution ($\mathbf{z}_{T}^{a}$) back to the clean data distribution ($\mathbf{z}_{0}$). Notably, the starting point of this trajectory contains adversarial noise, whereas the endpoint does not. Therefore, we aim to find a consistency model $\boldsymbol{f_{\theta}}$ that satisfies: $\boldsymbol{f_{\theta}}(\mathbf{z}_{t}^{a},\varnothing,t)=\boldsymbol{f_{\theta}}(\mathbf{z}_{t},\varnothing,t)=\mathbf{z}_{\epsilon}$, where $\mathbf{z}_{\epsilon}\approx\mathbf{z}_{0}$ denotes the limiting state of $\mathbf{z}_{t}$ as $t\to 0$. However, this contradicts Equation (7), as the trajectories initiated from $\mathbf{z}_{t}$ and $\mathbf{z}_{t}^{a}$ are misaligned, causing $\boldsymbol{f_{\theta}}(\mathbf{z}_{t}^{a},\varnothing,t)-\boldsymbol{f_{\theta}}(\mathbf{z}_{t},\varnothing,t)\to\boldsymbol{\epsilon}_{a}$. To explicitly reconcile this discrepancy, we introduce a coefficient $k_{t}$ and define an adjusted latent variable $\tilde{\mathbf{x}}_{t}$ to align the trajectories accordingly:
$\tilde{\mathbf{z}}_{t}=\mathbf{z}_{t}^{a}-k_{t}\boldsymbol{\epsilon}_{a},$ (9)
@@
with $k_{0}=1$ and $k_{T}=0$. Our goal is to ensure that the sampling distribution during the denoising process is independent of the adversarial perturbation $\boldsymbol{\epsilon}_{a}$. Although $\boldsymbol{\epsilon}_{a}$ can be computed during training as $\boldsymbol{\epsilon}_{a}=\arg\max_{\epsilon}\mathcal{L}(\mathrm{C}(\mathcal{D}(\mathbf{z}+\boldsymbol{\epsilon})),y_{\mathrm{true}})$, its exact value is unknown at inference time. Leveraging Bayes’ theorem and the properties of Gaussian distributions, we achieve this by selecting the value of coefficient $k_{t}$ such that the term involving $\boldsymbol{\epsilon}_{a}$ is eliminated. After a series of derivations, we obtain an explicit closed-form expression for $k_{t}$:
$k_{t}=\sqrt{\bar{\alpha}_{t}}-\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t})}{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{T})},0\leq t\leq T.$ (10)
Thus the $\tilde{\mathbf{z}}_{t}$ is constructed as:
$\tilde{\mathbf{z}}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}+\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t})}{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{T})}\boldsymbol{\epsilon}_{a}.$ (11)
@@
The full proof is provided in the Appendix.
@@
Accordingly, based on the loss function introduced in LCM (Luo et al. 2023a), our consistency distillation loss can be formulated as:
@@
$\begin{array}[]{l}\mathcal{L}_{CD}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{-}\right)=\\ \quad\mathbb{E}_{\mathbf{z},n}\left[d\left(\boldsymbol{f_{\theta}}\left(\tilde{\mathbf{z}}_{t_{n+k}},\varnothing,t_{n+k}\right),\boldsymbol{f_{\theta^{-}}}\left(\hat{\mathbf{z}}_{t_{n}}^{\Psi},\varnothing,t_{n}\right)\right)\right],\end{array}$ (12)
@@
where $d(\cdot,\cdot)$ denotes a distance metric, and $\Psi(\cdot,\cdot,\cdot,\cdot)$ represents the DDIM (Song, Meng, and Ermon 2022) PF ODE solver $\Psi_{\mathrm{DDIM}}$. The term $\hat{\mathbf{z}}_{t_{n}}^{\Psi}$ refers to the solution estimated by the solver when integrating from $t_{n+k}$ to $t_{n}$:
@@
Following (Kim et al. 2024), we also incorporate a reconstruction-like loss that leverages clean images to better align the distillation training process with the purification objective:
$\mathcal{L}_{\mathrm{rec}}\left(\boldsymbol{\theta}\right)=d\left(\boldsymbol{f_{\theta}}\left(\tilde{\mathbf{z}}_{t},\varnothing,t\right),\mathbf{z}\right)$ (14)
@@
To enhance the dynamical interpretability of the sampling process, we refine the DDIM-based PF ODE solver using a leapfrog-inspired mechanism. Specifically, we decompose the prediction into a position-like estimate of the clean image and a velocity-like estimate of the noise, which are then updated jointly through a first-order leapfrog integration step (Verlet 1967):
@@
where $\mathbf{z}_{0}=\sqrt{\bar{\alpha}_{t-1}}\cdot\hat{\mathbf{z}}_{0}$ represents the predicted clean image, while $\mathbf{v}_{0}=\sqrt{1-\bar{\alpha}_{t-1}}\cdot\hat{\boldsymbol{\epsilon}}$ serves as the midpoint velocity estimate.
Adaptive Semantic Enhanced Purification
Architecture Type Method Clean Acc. Robuse Acc.
$\ell_{\infty}$ $\ell_{1}$ $\ell_{2}$ Avg.
WRN-70-16 -- Vanilla 96.36 0.00 0.00 0.00 0.00
WRN-70-16 AT (Gowal et al. 2021a) 91.10 65.92 8.26 27.56 33.91
WRN-70-16 (Rebuffi et al. 2021) 88.54 64.26 12.06 32.29 36.20
WRN-70-16 Augment w/ Diff (Gowal et al. 2021b) 88.74 66.18 9.76 28.73 34.89
WRN-70-16 Augment w/ Diff (Wang et al. 2023) 93.25 70.72 8.48 28.98 36.06
MLP+WRN-28-10 AP (Shi, Holtz, and Mishne 2021) 91.89 4.56 8.68 7.25 6.83
UNet+WRN-70-16 (Yoon, Hwang, and Lee 2021) 87.93 37.65 36.87 57.81 44.11
UNet+WRN-70-16 GDMP (Wang et al. 2022) 93.16 22.07 28.71 35.74 28.84
UNet+WRN-70-16 ScoreOpt (Zhang et al. 2023a) 91.41 13.28 10.94 28.91 17.71
UNet+WRN-70-16 Purify++ (Zhang et al. 2023b) 92.18 43.75 39.84 55.47 46.35
UNet+WRN-70-16 DiffPure (Nie et al. 2022) 92.50 42.20 44.30 60.80 49.10
UNet+WRN-70-16 ADBM (Li et al. 2025) 91.90 47.70 49.60 63.30 53.50
UNet+WRN-70-16 DBLP (Ours) 94.8 58.4 59.4 64.4 60.73
[…]
Although diffusion models are effective at learning the denoising process from noise to images, relying solely on this process often leads to the loss of fine-grained details (Berrada et al. 2025). While OSCP (Lei et al. 2025) attempts to mitigate this by incorporating edge maps to enhance structural information, it uses fixed-threshold Canny edge detection (Canny 1986), which lacks adaptability to varying attack intensities. Moreover, adversarial perturbations introduce noise that can interfere with accurate edge extraction. To address these issues, we propose Adaptive Semantic Enhancement, a non-trainable, computationally efficient module to aggregate multi-scale edge information, enhancing structural integrity and detail preservation.

### v2

@@ 4.2 Noise Bridge Distillation
0: Dataset $\mathcal{D}$, LCM $\bm{f_{\theta}}$ and its initial model parameter $\bm{\theta}$, classifier $\mathrm{C}$, ground truth label $y_{\mathrm{true}}$, Leapfrog ODE solver $\Psi_{\mathrm{leap}}$, distance metric $d(\cdot,\cdot)$, EMA rate $\mu$, noise schedule $\alpha_{t}$, skip interval $k$, encoder $\mathcal{E}$; $\bm{\theta}^{-}\leftarrow\bm{\theta}$;
@@
4:  $\bm{\epsilon}_{a}=\arg\max_{\epsilon}\mathcal{L}(\mathrm{C}(\mathcal{D}(\mathbf{z}+\bm{\epsilon})),y_{\mathrm{true}})$;
5:  $\tilde{\mathbf{z}}_{t_{n+k}}=\sqrt{\bar{\alpha}_{t_{n+k}}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t_{n+k}}}\bm{\epsilon}+\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t_{n+k}})}{\sqrt{\bar{\alpha}_{t_{n+k}}}(1-\bar{\alpha}_{T})}\bm{\epsilon}_{a}$;
6:  ${\hat{\mathbf{z}}_{t_{n}}}^{\Psi_{\mathrm{leap}}}\leftarrow\mathbf{z}_{t_{n+k}}+\Psi(\mathbf{z}_{t_{n+k}},t_{n+k},t_{n},\varnothing)$;
@@
7:  $\mathcal{L}_{\mathrm{CD}}\left(\bm{\theta},\bm{\theta}^{-}\right)=d\left(\bm{f_{\theta}}\left(\tilde{\mathbf{z}}_{t_{n+k}},\varnothing,t_{n+k}\right),\bm{f_{\theta^{-}}}\left({\hat{\mathbf{z}}_{t_{n}}}^{\Psi_{\mathrm{leap}}},\varnothing,t_{n}\right)\right)$;
@@
8:  $\mathcal{L}_{\mathrm{rec}}\left(\bm{\theta}\right)=d\left(\bm{f_{\theta}}\left(\tilde{\mathbf{z}}_{t},\varnothing,t\right),\mathbf{z}\right)$;
9:  $\mathcal{L}\left(\bm{\theta},\bm{\theta}^{-}\right)=\mathcal{L}_{\mathrm{CD}}\left(\bm{\theta},\bm{\theta}^{-}\right)+\lambda_{\mathrm{rec}}\mathcal{L}_{\mathrm{rec}}\left(\bm{\theta}\right)$;
10:  $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\theta}\mathcal{L}(\bm{\theta},\bm{\theta}^{-})$;
11:  $\bm{\theta}^{-}\leftarrow\texttt{sg}(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta})$
12:end while
@@
Following LCM Luo et al. (2023a), let $\mathcal{E}$ and $\mathcal{D}$ denote the encoder and decoder that map images to and from the latent space, respectively. Given an image $\mathbf{x}$, its latent representation is $\mathbf{z}=\mathcal{E}(\mathbf{x})$. Unlike DDPM, DBLP includes adversarial perturbations $\bm{\epsilon}_{a}$ at the start of the noising process, such that $\mathbf{z}_{0}^{a}=\mathbf{z}_{0}+\bm{\epsilon}_{a}$. The forward process is then $\mathbf{x}_{t}^{a}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@
Our objective is to learn a trajectory that maps the diffused adversarial distribution ($\mathbf{z}_{T}^{a}$) back to the clean data distribution ($\mathbf{z}_{0}$). Notably, the starting point of this trajectory contains adversarial noise, whereas the endpoint does not. Therefore, we aim to find a consistency model $\bm{f_{\theta}}$ that satisfies: $\bm{f_{\theta}}(\mathbf{z}_{t}^{a},\varnothing,t)=\bm{f_{\theta}}(\mathbf{z}_{t},\varnothing,t)=\mathbf{z}_{\epsilon}$, where $\mathbf{z}_{\epsilon}\approx\mathbf{z}_{0}$ denotes the limiting state of $\mathbf{z}_{t}$ as $t\to 0$. However, this contradicts Equation equation 7, as the trajectories initiated from $\mathbf{z}_{t}$ and $\mathbf{z}_{t}^{a}$ are misaligned, causing $\bm{f_{\theta}}(\mathbf{z}_{t}^{a},\varnothing,t)-\bm{f_{\theta}}(\mathbf{z}_{t},\varnothing,t)\to\bm{\epsilon}_{a}$. To explicitly reconcile this discrepancy, we introduce a coefficient $k_{t}$ and define an adjusted latent variable $\tilde{\mathbf{x}}_{t}$ to align the trajectories accordingly:
$\tilde{\mathbf{z}}_{t}=\mathbf{z}_{t}^{a}-k_{t}\bm{\epsilon}_{a},$ (9)
@@
with $k_{0}=1$ and $k_{T}=0$. Our goal is to ensure that the sampling distribution during the denoising process is independent of the adversarial perturbation $\bm{\epsilon}_{a}$. Although $\bm{\epsilon}_{a}$ can be computed during training as $\bm{\epsilon}_{a}=\arg\max_{\epsilon}\mathcal{L}(\mathrm{C}(\mathcal{D}(\mathbf{z}+\bm{\epsilon})),y_{\mathrm{true}})$, its exact value is unknown at inference time. Leveraging Bayes’ theorem and the properties of Gaussian distributions, we achieve this by selecting the value of coefficient $k_{t}$ such that the term involving $\bm{\epsilon}_{a}$ is eliminated. After a series of derivations, we obtain an explicit closed-form expression for $k_{t}$:
$k_{t}=\sqrt{\bar{\alpha}_{t}}-\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t})}{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{T})},0\leq t\leq T,$ (10)
which satisfies $k_{0}=1$ and $k_{T}=0$. Thus the $\tilde{\mathbf{z}}_{t}$ is constructed as:
$\tilde{\mathbf{z}}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}+\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t})}{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{T})}\bm{\epsilon}_{a},$ (11)
@@
In this way, the sampling process doesn’t require $\bm{\epsilon}_{a}$. The full proof is provided in Appendix A.2.
Table 1: Clean Accuracy and Robust Accuracy ($\%$) results on CIFAR-10. Avg. denotes the average robust accuracy across three types of attack threats, vanilla refers to models without any adversarial defense mechanism. The best results are bolded, and the second best results are underlined.
Architecture Type Method
Clean
Acc.
Robuse Acc.
$\ell_{\infty}$ $\ell_{1}$ $\ell_{2}$ Avg.
WRN-70-16 - Vanilla 96.36 0.00 0.00 0.00 0.00
WRN-70-16 AT Gowal et al. (2021a) 91.10 65.92 8.26 27.56 33.91
WRN-70-16 Rebuffi et al. (2021) 88.54 64.26 12.06 32.29 36.20
WRN-70-16 Aug. w/ Diff Gowal et al. (2021b) 88.74 66.18 9.76 28.73 34.89
WRN-70-16 Aug. w/ Diff Wang et al. (2023) 93.25 70.72 8.48 28.98 36.06
MLP+WRN-28-10 AP Shi et al. (2021) 91.89 4.56 8.68 7.25 6.83
UNet+WRN-70-16 Yoon et al. (2021) 87.93 37.65 36.87 57.81 44.11
UNet+WRN-70-16 GDMP Wang et al. (2022) 93.16 22.07 28.71 35.74 28.84
UNet+WRN-70-16 ScoreOpt Zhang et al. (2023a) 91.41 13.28 10.94 28.91 17.71
UNet+WRN-70-16 Purify++ Zhang et al. (2023b) 92.18 43.75 39.84 55.47 46.35
UNet+WRN-70-16 DiffPure Nie et al. (2022) 92.50 42.20 44.30 60.80 49.10
UNet+WRN-70-16 ADBM Li et al. (2025) 91.90 47.70 49.60 63.30 53.50
UNet+WRN-70-16 DBLP (Ours) 94.8 58.4 64.4 59.4 60.73
@@
Accordingly, based on the loss function introduced in LCM Luo et al. (2023a), our consistency distillation loss can be formulated as:
@@
$\mathcal{L}_{CD}\left(\bm{\theta},\bm{\theta}^{-}\right)=\mathbb{E}_{\mathbf{z},n}\left[d\left(\bm{f_{\theta}}\left(\tilde{\mathbf{z}}_{t_{n+k}},\varnothing,t_{n+k}\right),\bm{f_{\theta^{-}}}\left(\hat{\mathbf{z}}_{t_{n}}^{\Psi},\varnothing,t_{n}\right)\right)\right],$ (12)
@@
where $d(\cdot,\cdot)$ denotes a distance metric, and $\Psi(\cdot,\cdot,\cdot,\cdot)$ represents the DDIM Song et al. (2022) PF ODE solver $\Psi_{\mathrm{DDIM}}$. The term $\hat{\mathbf{z}}_{t_{n}}^{\Psi}$ refers to the solution estimated by the solver when integrating from $t_{n+k}$ to $t_{n}$:
@@
Following Kim et al. (2024), we also incorporate a reconstruction-like loss that leverages clean images to better align the distillation training process with the purification objective:
$\mathcal{L}_{\mathrm{rec}}\left(\bm{\theta}\right)=d\left(\bm{f_{\theta}}\left(\tilde{\mathbf{z}}_{t},\varnothing,t\right),\mathbf{z}\right)$ (14)
@@
To enhance the dynamical interpretability of the sampling process, we refine the DDIM-based PF ODE solver using a leapfrog-inspired mechanism. Specifically, we decompose the prediction into a position-like estimate of the clean image and a velocity-like estimate of the noise, which are then updated jointly through a first-order leapfrog integration step Verlet (1967):
@@
where $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t-1}}\cdot\hat{\mathbf{z}}_{0}$ and $\mathbf{v}_{0}=\sqrt{1-\bar{\alpha}_{t-1}}\cdot\hat{\bm{\epsilon}}$, while $\mathbf{v}_{1/2}=2\mathbf{v}_{0}$ serves as the midpoint velocity estimate.
4.3 Adaptive Semantic Enhanced Purification
Table 2: Clean Accuracy and Robust Accuracy ($\%$) results on ImageNet. The default setting for attack is $\epsilon=4/255$. The best results are bolded, and the second best results are underlined.
Method Type Attack Standard Acc. Robust Acc. Architecture
w/o Defense - PGD-100 80.55 0.01 Res-50
Schott et al. (2019) AT PGD-40 72.70 47.00 Res-152
Wang et al. (2020) PGD-100 53.83 28.04 Res-50
ConvStem Singh et al. (2023) AutoAttack 77.00 57.70 ConvNeXt-L
MeanSparse Amini et al. (2024) AutoAttack 77.96 59.64 ConvNeXt-L
DiffPure Nie et al. (2022) AP PGD-100 68.22 42.88 Res-50
DiffPure Nie et al. (2022) AutoAttack 71.16 44.39 WRN-50-2
Bai et al. (2024) PGD-200 ($\epsilon=8/255$) 70.41 41.70 Res-50
Lee & Kim (2023) PGD+EOT 70.74 42.15 Res-50
Lin et al. (2025) PGD+EOT 68.75 45.90 Res-50
Zollicoffer et al. (2025) PGD-200 ($\epsilon=8/255$) 73.98 56.54 Res-50
MimicDiffusion Song et al. (2024) AutoAttack 66.92 61.53 Res-50
ScoreOpt Zhang et al. (2023a) Transfer-PGD 71.68 62.10 WRN-50-2
Pei et al. (2025) PGD-200 ($\epsilon=8/255$) 77.15 65.04 Res-50
OSCP Lei et al. (2025) PGD-100 77.63 73.89 Res-50
DBLP (Ours) PGD-100 78.2 75.6 Res-50
DBLP (Ours) AutoAttack 78.0 74.8 Res-50
DBLP (Ours) PGD-200 ($\epsilon=8/255$) 77.4 74.2 Res-50
Although diffusion models are effective at learning the denoising process from noise to images, relying solely on this process often leads to the loss of fine-grained details Berrada et al. (2025). While OSCP Lei et al. (2025) attempts to mitigate this by incorporating edge maps to enhance structural information, it uses fixed-threshold Canny edge detection Canny (1986), which lacks adaptability to varying attack intensities. Moreover, adversarial perturbations introduce noise that can interfere with accurate edge extraction. To address these issues, we propose Adaptive Semantic Enhancement, a non-trainable, computationally efficient module to aggregate multi-scale edge information, enhancing structural integrity and detail preservation.